# Production of Dark Photons through Higher Electromagnetic Moments at LDMX: Simulations and Model Discrimination

### v1

@@ 1 Introduction
The well established problem of the missing mass of the universe, based on undeniable evidence, has ignited the search for particles that can explain it. In more recent years, these Dark Matter (DM) searches have also been targeting masses outside of the canonical and widely explored Weakly Interacting Massive Particle (WIMP) [1, 2] mass window. The motivation to additionally look beyond WIMPs is as follows [3]: The lack of discovery and ever shrinking available parameter space of WIMPs could be explained by it being too light, lighter than the mass of a nucleon, to induce a sizeable nuclear recoil at direct detection experiments. If we take that this lighter than WIMP DM particle is thermal – was once in thermal equilibrium with the thermal bath – a new mediator particle is required due to the Lee-Weinberg bound [4]. There are several proposed mediators that provide a channel between Standard Model (SM) fields and DM. Among them the dark photon, the gauge field of a hypothetical new $U(1)$ symmetry group, is the most widely studied [5, 6, 7, 8, 9]. Hence, searches for this new mediator are complimentary to direct searches for probing these DM candidates. See [10] for a review on sub-GeV DM with a dark photon mediator.
@@
Fixed target experiments [11, 12, 13], such as the Light Dark Matter eXperiment (LDMX) [14, 15] and others [16, 17, 18, 19], focus on $\,\mathrm{MeV}-\,\mathrm{GeV}$ mass DM, so-called sub-GeV DM. At these experiments, a relativistic electron or proton beam is incident on a fixed target giving rise to interactions between the target nuclei and the beam particles. Depending on the experiment, there are various ways a DM signal is measured such as in a downstream DM detector or by measuring the missing energy/momentum of the recoil electrons. For the purpose of this study, we consider the feasibility of the latter case corresponding to the upcoming LDMX. Final state electrons which have significant missing energy and have gained transverse momentum (above expected background) are considered signal events. In this work, we are interested in signal events where a process analogous to ordinary photon bremsstrahlung occurs, dark bremsstrahlung [20]. Existing studies consider the production of dark photons [21] produced by dark bremsstrahlung, which decay invisibly into DM of various different models [22, 7, 23].
@@
The interaction between dark photons and SM fields emerge from vacuum polarization diagrams containing portal matter field loops [24] – leading to the ordinary kinetic mixing electromagnetic type interaction term. As described in [23], it is possible that the dark photon does not only couple through electromagnetic-like interactions, but also through higher order electromagnetic moment interactions. These additional interactions arise from loop diagrams with portal matter fields and new massive gauge bosons from a new gauge sector [25, 26].
@@
Due to the different Lorentz structure of these interactions between the dark photon and the SM fermions, the expected number of events and kinematic distributions of these events vary from those of the electromagnetic type interaction. This opens the opportunity for potential model differentiation in the case of an LDMX signal, and an extension of the projected sensitivity to include these alternative models. In this study, we extend the work of [23] to investigate the signatures and projected sensitivity of dark photons with kinetic mixing, magnetic and electric dipole, charge radius, and anapole moment interactions with SM fermions. Taking DM to be a complex scalar field as a benchmark, which naturally evades strong limits from indirect detection and energy injection into the cosmic microwave background [27, 28], we calculate the thermal targets of these models in order to enforce the constraints from Planck [29], putting the sensitivities into context. The analytic and numerical (using the Monte Carlo event generator MadGraph [30]) approach to calculating the kinematic distributions are compared, where we find that the common analytic method over estimates the number of high transverse momentum events. We discuss the plausibility of model and dark photon mass differentiation in the optimistic case of enough signal events at LDMX. Introducing a new variable, the model, which the kinematic distributions are dependent on, invokes a more difficult mass estimation. However, even though we find a degeneracy between mass and model in the kinematic distributions, considering both energy and transverse momentum of the recoil electron breaks the degeneracy between certain models.
@@ 2 Theoretical framework of Dark Moments
The effective interaction vertex between spin-1/2 fields and a photon can be parameterized in terms of five electromagnetic form factors, consisting of all possible Lorentz invariant operators [31]. These are effective interactions of higher order describing loop diagrams, inversely proportional to cut off scales of order of the loop field masses. The non-relativistic limit of these interactions reveal the static electromagnetic properties of the spin-1/2 field. The well studied kinetic mixing scenario of two $U(1)$ gauge groups [32, 33, 34], occurs through vacuum polarization diagrams containing new heavy portal matter fields carrying charges of both gauge groups. After integrating out the portal matter fields, we arrive at the effective kinetic mixing Lagrangian commonly written down. As shown in Ref. [24, 23], an analogous mechanism can generate new kinds of effective interactions beyond the common kinetic mixing scenario, including higher order electromagnetic moment interactions. Such higher order dark electromagnetic moment interactions (also known as form factors) occurs similarly in the standard model [35, 36, 37, 38, 39], and could occur in the dark sector between DM and the ordinary photon [40, 41, 42, 43, 44].
@@
We consider the scenario in which the SM is extended by a new $U(1)_{D}$ gauge group, with the associated gauge boson $A^{\prime}$, the dark photon [5], which receives a mass via the Stueckelberg mechanism or a dark Higgs mechanism. The existence of a new set of fields, portal matter fields [^1] It has been shown in [24] that the portal matter fields are likely fermions, and if so must be vector-like with respect to the SM and dark gauge groups., carrying SM and dark $U(1)_{D}$ charges allows for the interaction between the $A^{\prime}$ and SM fermions through vacuum polarization diagrams with these portal matter fields in the loops [24]. Previous literature has discussed examples of UV complete frameworks which contain these heavy portal matter fields [25, 24, 26, 45]. As discussed in [23, 46], these loop diagrams generate additional interactions to the standard electromagnetic type interaction between $A^{\prime}$ and SM fermions $f$. We consider in this study four additional types of interactions, separately [^2] A scenario where $\epsilon$ is much smaller than the other effective interaction couplings is possible, since $\epsilon$ depends on the extended Higgs sector generating the mass of the dark photon [23]. Therefore we justifiably consider the implications for scenarios where each interaction is considered in isolation.. We are justified to write these interactions in an EFT framework as higher dimensional interactions due to the small enough energy scale in which we are interested in. In general the strengths of these interactions are dependent on momentum and a cut off scale $\Lambda_{f}$. As described in [23] and calculated in [47], $\Lambda_{f}$ is dependent on the portal matter field mass. We take the mass of the portal matter to be at or above the TeV scale to evade LHC constraints, leading to $\Lambda_{f}\sim\mathcal{O}(\,\mathrm{TeV})$. We safely neglect these dependencies since we are interested in sub-GeV masses, thus energies far enough below the TeV scale. We write these interactions as follows,
@@
$\mathscr{L}_{A^{\prime}f}=$ $-e\epsilon Q_{f}\overline{f}\gamma^{\mu}fA^{\prime}_{\mu}-\mu_{f}\partial_{\nu}\left(\overline{f}\sigma^{\mu\nu}f\right)A^{\prime}_{\mu}-id_{f}\partial_{\nu}\left(\overline{f}\sigma^{\mu\nu}\gamma^{5}f\right)A^{\prime}_{\mu}$ (2.1)
@@
where $f$ are the SM fermions and $e$ is the elementary charge. The first term is the usual kinetic mixing term with interaction strength proportional to $\epsilon$, in addition to the magnetic dipole term with coupling $\mu_{f}$, electric dipole with $d_{f}$, charge radius with $b_{f}$, and anapole moment with $a_{f}$, which we take to be free parameters. We define the parameter $g_{f}$ representing the coupling strength for each type of interaction, or model,
$g_{f}=\begin{cases}\epsilon&\text{Kinetic Mixing}\\ \mu_{f}/e&\text{Magnetic Dipole}\\ d_{f}/e&\text{Electric Dipole}\\ b_{f}/e&\text{Charge Radius}\\ a_{f}/e&\text{Anapole Moment}\\ \end{cases}$ (2.2)
where we compare the phenomenology of each. Notice that, for all cases except for kinetic mixing, $g_{f}$ is dimensionful. The magnetic and electric dipole models are expected to have similar phenomenology, since their Lorentz structures only differ by a $\gamma^{5}$ thus their cross sections scale the same way with energy; similarly with the charge radius and anapole moment models. This allows for the models to be classified into three groups: the first with the kinetic mixing, the second with the magnetic and electric dipole, and the third with the charge radius and anapole moment models.
@@
$\mathscr{L}_{DM}=|\partial_{\mu}\chi|^{2}-m_{\text{DM}}^{2}|\chi|^{2}+ig_{\text{D}}A^{\prime\mu}\left[\chi^{*}\left(\partial_{\mu}\chi\right)-\left(\partial_{\mu}\chi^{*}\right)\chi\right],$ (2.3)
where $m_{\text{DM}}$ is $O(\,\mathrm{MeV}-\,\mathrm{GeV})$, $m_{A^{\prime}}/m_{\text{DM}}=3$, and we quantify in terms of the dark fine structure constant, $\alpha_{D}\equiv\frac{g_{D}^{2}}{4\pi}$. We take $\alpha_{D}=0.5$, and $g_{D}\gg g_{f}$ such that $\text{BR}(A^{\prime}\to\chi\chi)\approx 1$. Due to its p-wave dominant thermally averaged annihilation cross section, the scalar DM candidate is safe from strong indirect detection constraints [48].
@@ 3 Relic Density
We enforce consistency with the Planck measurement of the DM relic abundance [29],
$\Omega_{\text{DM,obs}}h^{2}=0.120\pm 0.001,$ (3.1)
@@
by solving the Boltzmann equation [49], which calculates the relic abundance as a function of model parameters for each model,
$\dot{n}_{\chi}+3Hn_{\chi}=-\frac{1}{2}\langle\sigma v_{\rm rel}\rangle_{\chi\overline{\chi}\rightarrow f{\overline{f}}}\,(n_{\chi}^{2}-n_{\chi,{\rm eq}}^{2}),$ (3.2)
@@
where $n_{\chi}$ is the cosmological number density of DM particles, $H$ is the Hubble rate, and $\langle\sigma v_{\rm rel}\rangle_{\chi\overline{\chi}\to f\overline{f}}$ is the thermally averaged cross section of DM annihilation into fermions. $\langle\sigma v_{\rm rel}\rangle$ is calculated using the matrix elements presented in Appendix A.
@@
We use our own Boltzmann solver based on [49], which is validated with micrOMEGAs[50] version 5 and includes contributions from DM annihilation into hadronic final states [51, 52]. For the sake of this study, we consider large enough couplings, $\alpha_{D}=0.5$, that DM reaches thermal equilibrium with the SM bath therefore the abundance is set through the freeze-out mechanism. Due to our choice of $m_{A^{\prime}}/m_{\text{DM}}=3$, DM annihilation into SM fermions through an s-channel with a virtual dark photon is the dominant process setting the abundance.
@@
Figure 1: Projected sensitivity of LDMX [14] of each model with their relic targets in the $y$ vs $m_{A^{\prime}}$ plane, where $g_{f}$ is defined in Section 2. We take the set up of Phase II LDMX, which has $10^{16}$ EOT and an $8\,\mathrm{GeV}$ beam, for the optimistic zero background scenario at 90% confidence level. We set $m_{A^{\prime}}/m_{\text{DM}}=3$, and $\alpha_{D}=0.5$. The kinetic mixing (KM) is shown in red, magnetic dipole (M) in green, electric dipole (E) in blue, charge radius (C) in grey, and anapole moment (A) in brown, while the LDMX sensitivity is plotted with dashed curves and the relic targets with solid curves.
@@
Fig. 1 plots the relic targets, the contours in $y=g_{f}^{2}\alpha_{D}\left(\frac{m_{\text{DM}}}{m_{A^{\prime}}}\right)^{4}$ vs $m_{A^{\prime}}$ space corresponding to the observed relic abundance, for each model in tandem with the projected exclusion bounds described in the following section. As described in the previous section, the models can be separated into three groups, each leading to quantitatively similar relic targets: the first with the kinetic mixing, the second with the magnetic and electric dipole, and the third with the charge radius and anapole moment models. By examining the behaviour of the squared amplitudes of the process dominantly setting the relic density, found in Appendix A, this will tell us about the thermally averaged cross section scaling with $m_{A^{\prime}}$ near freeze-out. The relic target of the kinetic mixing model is validated with previous results [22], with its positive slope. The magnetic and electric dipole models exhibit an approximately constant relic target as a function of dark photon mass. As discussed in [23] for the magnetic dipole model, which can also be applied to the electric dipole model, the energy-squared proportionality of the DM annihilation matrix element squared leads to a cross section near freeze-out that is approximately $m_{A^{\prime}}$ independent. Finally, the charge radius and anapole moment model group’s relic target slope is due to the additional energy squared proportionality in the matrix element squared, resulting in a cross section near freeze-out proportional to $m_{A^{\prime}}^{2}$.

### v2

@@ 1 Introduction
The well established problem of the missing mass of the universe, based on undeniable evidence, has ignited the search for particles that can explain it. In more recent years, these Dark Matter (DM) searches have also been targeting masses outside of the canonical and widely explored Weakly Interacting Massive Particle (WIMP) Arcadi:2017kky ; Roszkowski:2017nbc mass window. The motivation to additionally look beyond WIMPs is as follows Essig:2011nj : The lack of discovery and ever shrinking available parameter space of WIMPs could be explained by it being too light, lighter than the mass of a nucleon, to induce a sizeable nuclear recoil at direct detection experiments. If we take that this lighter than WIMP DM particle is thermal – was once in thermal equilibrium with the thermal bath – a new mediator particle is required due to the Lee-Weinberg bound Lee:1977ua . There are several proposed mediators that provide a channel between Standard Model (SM) fields and DM. Among them the dark photon, the gauge field of a hypothetical new $U(1)$ symmetry group, is the most widely studied Fabbrichesi:2020wbt ; Cosme:2021baj ; Catena:2023use ; Curtin:2014cca ; Graham:2021ggy . Hence, searches for this new mediator are complimentary to direct searches for probing these DM candidates. See Balan:2024cmq for a review on sub-GeV DM with a dark photon mediator.
@@
Fixed target experiments Antel:2023hkf ; Ilten:2022lfq ; Beacham:2019nyx , such as the Light Dark Matter eXperiment (LDMX) LDMX:2018cma ; Akesson:2022vza and others NA64:2023wbi ; Belle-II:2018jsg ; COHERENT:2021pvd ; SHiP:2020noy , focus on $\mathrm{MeV}-\,\mathrm{GeV}$ mass DM, so-called sub-GeV DM. At these experiments, a relativistic electron or proton beam is incident on a fixed target giving rise to interactions between the target nuclei and the beam particles. Depending on the experiment, there are various ways a DM signal is measured such as in a downstream DM detector or by measuring the missing energy/momentum of the recoil electrons. For the purpose of this study, we consider the feasibility of the latter case corresponding to the upcoming LDMX. Final state electrons which have significant missing energy and have gained transverse momentum (above expected background) are considered signal events. In this work, we are interested in signal events where a process analogous to ordinary photon bremsstrahlung occurs, dark bremsstrahlung Bjorken:2009mm . Existing studies consider the production of dark photons Cline:2024qzv produced by dark bremsstrahlung, which decay invisibly into DM of various different models Berlin:2018bsc ; Catena:2023use ; Rizzo:2021lob .
@@
The interaction between dark photons and SM fields emerge from vacuum polarization diagrams containing portal matter field loops Rizzo:2018vlb – leading to the ordinary kinetic mixing electromagnetic type interaction term. As described in Rizzo:2021lob ; Barducci:2021egn , it is possible that the dark photon does not only couple through electromagnetic-like interactions, but also through higher order electromagnetic moment interactions. These additional interactions arise from loop diagrams with portal matter fields and new massive gauge bosons from a new gauge sector Rueter:2019wdf ; Wojcik:2020wgm .
@@
Due to the different Lorentz structure of these interactions between the dark photon and the SM fermions, the expected number of events and kinematic distributions of these events vary from those of the electromagnetic type interaction. This opens the opportunity for potential model differentiation in the case of an LDMX signal, and an extension of the projected sensitivity to include these alternative models. This concept is also employed in Blinov:2020epi , where the authors consider the potential for model discrimination at missing momentum experiments. In this study, we extend the work of Rizzo:2021lob to investigate the signatures and projected sensitivity of dark photons with kinetic mixing, magnetic and electric dipole, charge radius, and anapole moment interactions with SM fermions. Taking DM to be a complex scalar field as a benchmark, which naturally evades strong limits from indirect detection and energy injection into the cosmic microwave background Cirelli:2023tnx ; Slatyer:2015kla , we calculate the thermal targets of these models in order to enforce the constraints from Planck Planck:2018vyg , putting the sensitivities into context. The analytic and numerical (using the Monte Carlo event generator MadGraph Alwall:2014hca ) approach to calculating the kinematic distributions are compared, where we find that the common analytic method over estimates the number of high transverse momentum events. We discuss the plausibility of model and dark photon mass differentiation in the optimistic case of enough signal events at LDMX. Introducing a new variable, the model, which the kinematic distributions are dependent on, invokes a more difficult mass estimation. However, even though we find a degeneracy between mass and model in the kinematic distributions, considering both energy and transverse momentum of the recoil electron breaks the degeneracy between certain models.
@@ 2 Theoretical framework of Dark Moments
The effective interaction vertex between spin-1/2 fields and a photon can be parameterized in terms of five electromagnetic form factors, consisting of all possible Lorentz invariant operators Nowakowski:2004cv . These are effective interactions of higher order describing loop diagrams, inversely proportional to cut off scales of order of the loop field masses. The non-relativistic limit of these interactions reveal the static electromagnetic properties of the spin-1/2 field. The well studied kinetic mixing scenario of two $U(1)$ gauge groups Holdom:1985ag ; Dienes:1996zr ; DelAguila:1993px , occurs through vacuum polarization diagrams containing new heavy portal matter fields carrying charges of both gauge groups. After integrating out the portal matter fields, we arrive at the effective kinetic mixing Lagrangian commonly written down. As shown in Ref. Rizzo:2018vlb ; Rizzo:2021lob , an analogous mechanism can generate new kinds of effective interactions beyond the common kinetic mixing scenario, including higher order electromagnetic moment interactions. Such higher order dark electromagnetic moment interactions (also known as form factors) occurs similarly in the standard model Dombey:1980eu ; Schwinger:1948iu ; Muong-2:2003xkd ; Hoogeveen:1990cb ; Bernstein:1963qh , and could occur in the dark sector between DM and the ordinary photon Barger:2010gv ; Alves:2017uls ; Lambiase:2021xcj ; Pospelov:2000bq ; Hambye:2021xvd .
@@
We consider the scenario in which the SM is extended by a new $U(1)_{D}$ gauge group, with the associated gauge boson $A^{\prime}$, the dark photon Fabbrichesi:2020wbt , which receives a mass via the Stueckelberg mechanism or a dark Higgs mechanism. The existence of a new set of fields, portal matter fields [^1] It has been shown in Rizzo:2018vlb that the portal matter fields are likely fermions, and if so must be vector-like with respect to the SM and dark gauge groups., carrying SM and dark $U(1)_{D}$ charges allows for the interaction between the $A^{\prime}$ and SM fermions through vacuum polarization diagrams with these portal matter fields in the loops Rizzo:2018vlb . Previous literature has discussed examples of UV complete frameworks which contain these heavy portal matter fields Rueter:2019wdf ; Rizzo:2018vlb ; Wojcik:2020wgm ; Rueter:2020qhf . As discussed in Rizzo:2021lob ; Chu:2018qrm , these loop diagrams generate additional interactions to the standard electromagnetic type interaction between $A^{\prime}$ and SM fermions $f$. We consider in this study four additional types of interactions, separately [^2] A scenario where $\epsilon$ is much smaller than the other effective interaction couplings is possible, since $\epsilon$ depends on the extended Higgs sector generating the mass of the dark photon Rizzo:2021lob . Therefore we justifiably consider the implications for scenarios where each interaction is considered in isolation.. We are justified to write these interactions in an EFT framework as higher dimensional interactions due to the small enough energy scale in which we are interested in. In general, the strengths of these interactions are dependent on the squared momentum of the $A^{\prime}$, labeled $q^{2}$, and a cut off scale, $\Lambda_{f}$, where we can write the form factors as an expansion around $q^{2}=0$ as $c_{0}+c_{1}q^{2}/\Lambda_{f}^{2}+\dots$, where $c_{i}$ are constants. As described in Rizzo:2021lob and calculated in Leveille:1977rc , $\Lambda_{f}$ is dependent on the portal matter field masses. Since the masses of the portal matter must be at or above the TeV scale to evade LHC constraints, we set $\Lambda_{f}=1\,\mathrm{TeV}$ for the purposes of this study. We safely neglect the higher dimensional terms in the form factor, beyond $c_{0}$, since we are interested in sub-GeV masses, thus energies/momenta far enough below the TeV scale. We write these interactions as follows,
@@
$\mathscr{L}_{A^{\prime}f}=$ $-e\epsilon Q_{f}\overline{f}\gamma^{\mu}fA^{\prime}_{\mu}-\mu_{f}\partial_{\nu}\left(\overline{f}\sigma^{\mu\nu}f\right)A^{\prime}_{\mu}-id_{f}\partial_{\nu}\left(\overline{f}\sigma^{\mu\nu}\gamma^{5}f\right)A^{\prime}_{\mu}$ (1)
@@
where $f$ are the SM fermions [^3] Coupling to neutrinos is assumed to be absent, and for the purposes of this study we mainly consider couplings to electrons and muons. and $e$ is the elementary charge. The first term is the usual kinetic mixing term with interaction strength proportional to $\epsilon$, in addition to the magnetic dipole term with coupling $\mu_{f}$, electric dipole with $d_{f}$, charge radius with $b_{f}$, and anapole moment with $a_{f}$, which we take to be free parameters. We define the parameter $g_{f}$ representing the coupling strength for each type of interaction, or model,
[…]
where we compare the phenomenology of each. Notice that, for all cases except for kinetic mixing, $g_{f}$ is dimensionful with dimension -1 for the magnetic and electric dipole, and -2 for the charge radius and anapole moment. The magnetic and electric dipole models are expected to have similar phenomenology, since their Lorentz structures only differ by a $\gamma^{5}$ thus their cross sections scale the same way with energy; similarly with the charge radius and anapole moment models. This allows for the models to be classified into three groups: the first with the kinetic mixing, the second with the magnetic and electric dipole, and the third with the charge radius and anapole moment models.
@@
$\mathscr{L}_{DM}=|\partial_{\mu}\chi|^{2}-m_{\text{DM}}^{2}|\chi|^{2}+ig_{\text{D}}A^{\prime\mu}\left[\chi^{*}\left(\partial_{\mu}\chi\right)-\left(\partial_{\mu}\chi^{*}\right)\chi\right],$ (3)
where $m_{\text{DM}}$ is $O(\,\mathrm{MeV}-\,\mathrm{GeV})$, $m_{A^{\prime}}/m_{\text{DM}}=3$, and we quantify in terms of the dark fine structure constant, $\alpha_{D}\equiv\frac{g_{D}^{2}}{4\pi}$. We take $\alpha_{D}=0.5$, and $g_{D}\gg g_{f}$ such that $\text{BR}(A^{\prime}\to\chi\chi)\approx 1$.
Conservation of angular momentum forbids scalar particle s-wave annihilation via an s-channel dark photon – giving rise to its p-wave dominant (velocity suppressed) thermally averaged annihilation cross section. This argument is independent from the type of interaction between the dark photon and SM fermions. Therefore, the scalar DM candidate is safe from strong indirect detection constraints in these models Boehm:2003hm ; Cirelli:2021fbo .
Direct detection of sub-GeV dark matter primarily relies on scattering with electrons or the excitation of vibrational modes in the detector material Essig:2024wtj . As demonstrated in Rizzo:2021lob for the magnetic moment model, the direct detection signal for electron recoils is expected to be small. We expect similar estimates for the electric dipole, charge radius, and anapole dark photon models. Although dedicated studies of the direct detection prospects for these higher electromagnetic moment dark photons present an intriguing direction for future study, we do not consider them in this work.
@@ 3 Relic Density
We enforce consistency with the Planck measurement of the DM relic abundance Planck:2018vyg ,
$\Omega_{\text{DM,obs}}h^{2}=0.120\pm 0.001,$ (4)
@@
by solving the Boltzmann equation Gondolo:1990dk , which calculates the relic abundance as a function of model parameters for each model,
$\dot{n}_{\chi}+3Hn_{\chi}=-\frac{1}{2}\langle\sigma v_{\rm rel}\rangle_{\chi\chi\rightarrow ff}\,(n_{\chi}^{2}-n_{\chi,{\rm eq}}^{2}),$ (5)
@@
where $n_{\chi}$ is the cosmological number density of DM particles, $H$ is the Hubble rate, and $\langle\sigma v_{\rm rel}\rangle_{\chi\chi\to ff}$ is the thermally averaged cross section of DM annihilation into fermions, namely electrons and muons. $\langle\sigma v_{\rm rel}\rangle$ is calculated using the matrix elements presented in Appendix A.
@@
We use our own Boltzmann solver based on Gondolo:1990dk , which is validated with micrOMEGAsBelanger:2013oya version 5. For the kinetic mixing model, we include contributions from DM annihilation into hadronic final states PDG ; Ilten:2018crw ; Izaguirre:2015zva . For the sake of this study, we consider large enough couplings, $\alpha_{D}=0.5$, that DM reaches thermal equilibrium with the SM bath therefore the abundance is set through the freeze-out mechanism. Due to our choice of $m_{A^{\prime}}/m_{\text{DM}}=3$, DM annihilation into SM fermions through an s-channel with a virtual dark photon is the dominant process setting the abundance.
@@
Figure 1: Projected sensitivity of LDMX LDMX:2018cma of each model with their relic targets in the $y$ vs $m_{A^{\prime}}$ plane, where $g_{f}$ is defined in Section 2. We take the set up of Phase II LDMX, which has $10^{16}$ EOT and an $8\,\mathrm{GeV}$ beam, for the optimistic zero background scenario at 90% confidence level. We set $m_{A^{\prime}}/m_{\text{DM}}=3$, and $\alpha_{D}=0.5$. The kinetic mixing (KM) is shown in red, magnetic dipole (M) in green, electric dipole (E) in blue, charge radius (C) in grey, and anapole moment (A) in brown, while the LDMX sensitivity is plotted with dashed curves and the relic targets with solid curves.
Fig. 1 plots the relic targets, the contours corresponding to the observed relic abundance in $y$ vs $m_{A^{\prime}}$ space for each model in tandem with the projected exclusion bounds described in the following section. $y$ is defined to be,
$y\equiv g_{f}^{2}\alpha_{D}\left(\frac{m_{\text{DM}}}{m_{A^{\prime}}}\right)^{4}\Lambda_{f}^{2n},$ (6)
where $\alpha_{D}$ and $g_{f}$ are defined in Section 2, and $n=1$ for the electric and magnetic dipole, and n=2 for the anapole and charge radius. We take an adapted version of the commonly used $y$ parameter, a convenient choice since the early universe DM annihilation cross section $\times$ velocity, $\sigma v_{\chi\chi\to ff}\propto\frac{y}{m_{\text{DM}}^{2}}$ LDMX:2018cma . We include the factor $\Lambda_{f}^{2n}$ to have a dimensionless $y$ for all models. For the higher order models we include a shaded region to the relic targets in Fig. 1, in which we expect contributions for hadronic final states. However, we do not explicitly compute these contributions for models beyond kinetic mixing interactions [^4] The shaded regions in the relic targets are computed assuming Eq. A4 of Izaguirre:2015zva . This computation is not valid for non-vector interactions, and we leave the development of a method to better approximate other types of dark photon interactions for future work. These shaded regions are only to be used as a proxy for the contribution from hadronic final states – not to be taken rigorously..
@@
As described in the previous section, the models can be separated into three groups, each leading to quantitatively similar relic targets: the first with the kinetic mixing, the second with the magnetic and electric dipole, and the third with the charge radius and anapole moment models. By examining the squared amplitudes of the process dominantly setting the relic density, found in Appendix A, this will tell us about the thermally averaged cross section scaling with $m_{A^{\prime}}$ near freeze-out. The relic target of the kinetic mixing model is validated with previous results Berlin:2018bsc , with its positive slope. The magnetic and electric dipole models exhibit an approximately constant relic target as a function of dark photon mass. As discussed in Rizzo:2021lob for the magnetic dipole model, which can also be applied to the electric dipole model, the energy-squared proportionality of the DM annihilation matrix element squared leads to a cross section near freeze-out that is approximately $m_{A^{\prime}}$ independent. Finally, the charge radius and anapole moment model group’s relic target slope is due to the additional energy squared proportionality in the matrix element squared, resulting in a cross section near freeze-out proportional to $m_{A^{\prime}}^{2}$.